# A conjecture of Radu and Sellers on congruences modulo powers of 2 for broken 3-diamond partitions

### v1

@@ 4. Radu and Sellers’ conjecture
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $k=l~{}~{}\text{or}~{}~{}l+1$ there exist polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$ such that
@@
(4.2) $ord_{t}(p_{k}^{(i)}(t))\geq\lceil\frac{k+s_{i}}{2}\rceil,i\in\{1,2\},$
for some fixed integers $s_{1}$ and $s_{2}$. Then there exist families of polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$, integer $k\geq l$, such that (4.1) and (4.2) hold for integer $k\geq l$.
@@
Let $N\textgreater l+1$ be an integer and assume by induction that there are families of polynomials $p_{k}^{(i)}(t),i\in\{1,2\}$, such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
$p_{k}^{(i)}(t)=\sum_{n\geq\lceil\frac{k+s_{i}}{2}\rceil}c_{i}(k,n)t^{n},l\leq k\leq N-1,$
with integers $c_{i}(k,n)$. Apply Lemma 3.3 we get:
@@
$=$ $-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n\geq\lceil\frac{N+j-2+s_{i}}{2}\rceil}c_{i}(N+j-2,n)t^{n}$
$=$ $-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}a_{j}(t)t^{-1}\sum_{n\geq\lceil\frac{N+j+s_{i}}{2}\rceil}c_{i}(N+j-2,n-1)t^{n}.$
@@
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose for $k=l~{}~{}\text{or}~{}~{}l+1$ there exist polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$, $i\in\{1,2\}$, such that
@@
(4.4) $p_{k}^{(i)}(t)=\sum_{n}c_{i}(k,n)2^{\lfloor\frac{2n-k+\gamma_{i}}{2}\rfloor}t^{n},$
with integers $\gamma_{i}$ and $c_{i}(k,n)$. Then there exist families of polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$, integer $k\geq l$, of the form (4.4) for which property (4.3) holds for integer $k\geq l$.
@@
Suppose for an integer $N\textgreater l+1$ there are families of polynomials $p_{k}^{(i)}(t)$, $i\in\{1,2\}$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$. We proceed by mathematical induction on $N$. Applying Lemma 3.3 and using the induction base (4.3) and (4.4) we attain:
$U_{2}(ut^{N})=-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)2^{\lfloor{\frac{2n-(N+j-2)+\gamma_{i}}{2}}\rfloor}t^{n}.$
Utilizing (3.7) this rewrites into:
$U_{2}(ut^{N})$ $=-\sum_{j=0}^{1}\sum_{l=1}^{2}s(j,l)2^{\lfloor{\frac{2l+j-1}{2}}\rfloor}t^{l}\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)2^{\lfloor{\frac{2n-(N+j-2)+\gamma_{i}}{2}}\rfloor}t^{n}$
$=-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}\sum_{l=1}^{2}\sum_{n}s(j,l)c_{i}(N+j-2,n-l)2^{\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}}\rfloor+\lfloor{\frac{2l+j-1}{2}}\rfloor}t^{n}.$
@@
$\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}}\rfloor+\lfloor{\frac{2l+j-1}{2}}\rfloor$
$\geq\lfloor{\frac{2(n-l)-(N+j-2)+\gamma_{i}+2l+j-2}{2}}\rfloor$
$=\lfloor{\frac{2n-N+\gamma_{i}}{2}}\rfloor.$
@@
A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support. A map $a:\mathbb{Z}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called discrete array if for each $i\in\mathbb{Z}$ the map $a(i,-):\mathbb{Z}\rightarrow\mathbb{Z},j\mapsto a(i,j)$, has finite support.
@@
Given $U^{(0)}(f)$, $U^{(1)}(f)$ as in Definition 2.2, $p_{0},p_{1}$, and $y_{0}$ as in Definition 2.5. Then there exist discrete arrays $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$, $i\in\{0,1\}$, such that the following relations hold for all $k\in\mathbb{N^{\ast}}$:
(4.5) $U^{(0)}(t^{k})=p_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{0}(k,n)2^{\lfloor\frac{2n-k+2}{2}\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{1}(k,n)2^{\lfloor\frac{2n-k+2}{2}\rfloor}t^{n},$
(4.6) $U^{(0)}(p_{1}t^{k})=p_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{0}(k,n)2^{\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{1}(k,n)2^{\lfloor\frac{2n-k-1}{2}\rfloor}t^{n},$
(4.7) $U^{(1)}(p_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}c_{0}(k,n)2^{\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}c_{1}(k,n)2^{\lfloor\frac{2n-k-3}{2}\rfloor}t^{n},$
(4.8) $U^{(1)}(y_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{0}(k,n)2^{\lfloor\frac{2n-k}{2}\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{1}(k,n)2^{\lfloor\frac{2n-k-4}{2}\rfloor}t^{n}.$
@@
The Appendix lists eight fundamental relations. The two fundamental relations of Group I fit the pattern of the relation (4.5) for two values of $k$. The same observations applies to the Groups II, III and IV with regard to the relations (4.6), (4.8), and (4.7), respectively. Applying Lemmas 4.1 and 4.2 immediately proves the statement for all $k\in\mathbb{N^{*}}$. ∎
@@
by Garvan’s MAPLE package ETA (3.1). Then we note that $p$ and $p_{1}$ have expansions in powers of $q$ with coefficients in $\mathbb{Z}$. Since $(1-q^{k})^{2}\equiv 1-q^{2k}\pmod{2}$, we have $\eta_{k}^{2}\equiv\eta_{2k}\pmod{2}$. Then
@@
which implies that function $p$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$. We now have that the coefficients of $4p_{1}=p+2pt+t+4t^{2}$ have integral coefficients. Hence, by the relations of $U^{(1)}(p_{0}t^{2})$ given in Appendix,
@@
where $f(t)$ and $g(t)$ are polynomial of $t$ with integral coefficients, we see that $p_{1}t$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$ as well $p_{1}$. Furthermore, using (4.9) and combining the relations of $U^{(1)}(p_{0}t)$, $U^{(1)}(y_{0}t)$, $U^{(0)}(p_{1}t)$ given in Appendix and $U^{(0)}(t^{4})$ given in (3.3), we now state the following representation.
@@
$L_{3}=$ $y_{0}t+2y_{0}t^{2}+4g(p_{0},y_{0},t).$
for some polynomial $f$ and $g$ with integral coefficients.
@@
For $\alpha\geq 1$ there exist integers $d_{n,i}^{(\alpha)}$, $i=0,1,2,3$ such that
@@
By (4.9) we see that $L_{1}=-4p_{0}t+y_{0}t$ has form given in (4.11). Assume that $L_{2\alpha-1}$ has the form given in (4.11) for a fixed $\alpha$, then
@@
We aim to show that the form of each sum on the right-hand side of (4.13) satisfies the form in (4.12). From (4.7) and (4.8),
$L_{2\alpha}=$ $p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{0}(k,n)2^{k-1+\lfloor\frac{2n-k+1}{2}\rfloor}t^{n}+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{1}(k,n)2^{k-1+\lfloor\frac{2n-k-3}{2}\rfloor}t^{n}$
$+p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{0}(k,n)2^{k-1+\lfloor\frac{2n-k}{2}\rfloor}t^{n}+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{1}(k,n)2^{k-1+\lfloor\frac{2n-k-4}{2}\rfloor}t^{n}.$
@@
$k-1+\lfloor\frac{2n-k+1}{2}\rfloor=\lfloor\frac{2n-k+1+2k-2}{2}\rfloor=n+\lfloor\frac{k-1}{2}\rfloor\geq n,$
$k-1+\lfloor\frac{2n-k}{2}\rfloor=\lfloor\frac{2n-k+2k-2}{2}\rfloor=n+\lfloor\frac{k-2}{2}\rfloor\geq n-1.$
@@
$k-1+\lfloor\frac{2n-k-3}{2}\rfloor=\lfloor\frac{2n-k-3+2k-2}{2}\rfloor=n+\lfloor\frac{k-5}{2}\rfloor\geq n-2,$
$k-1+\lfloor\frac{2n-k-4}{2}\rfloor=\lfloor\frac{2n-k-4+2k-2}{2}\rfloor=n+\lfloor\frac{k-6}{2}\rfloor\geq n-3.$
So that the sums in the right-hand side of (4.13) has the form of (4.12). Hence $L_{2\alpha}$ has the desired form. The proof that the correct form of $L_{2\alpha}$ implies the correct form of $L_{2\alpha+1}$ is analogous. The general result follows by induction. ∎
By Lemma 4.2 and Lemma 4.6 we let
@@
and denote $\pi(n)$ be the 2-adic order of n (i.e. the highest power of 2 that divides $n$). By Lemma 4.6, we should notice that for $m,n\in\mathbb{N}$ and $\alpha\geq 1$, $l_{n,1}$ and $l_{m,3}$ only appear in $L_{2\alpha}$, while $l_{n,0}$ and $l_{m,2}$ only appear in $L_{2\alpha-1}$.
@@
    (4.14) $\pi(D^{(2\alpha-1)}(l_{m,i},l_{n,j}))\geq 2\alpha-2+m+n+\lambda_{i,j},$
    (4.15) $\pi(D^{(2\alpha-1)}(l_{1,2},l_{2,0}))\geq 2\alpha,$
    (4.16) $\pi(D^{(2\alpha-1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2\alpha;$
@@
    (4.17) $\pi(D^{(2\alpha)}(l_{m,i},l_{n,j}))\geq 2\alpha-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
    (4.18) $\pi(D^{(2\alpha)}(l_{1,1},l_{3,3}))\geq 2\alpha,$
    (4.19) $\pi(D^{(2\alpha)}(l_{2,1},l_{3,3}))\geq 2\alpha+1;$
@@
$\lambda_{i,j}=\left\{\begin{array}[]{ll}-2,&\textrm{if $(i,j)=(1,1)$, $(0,2)$, $(2,0)$, or $(0,0)$},\\ -3,&\textrm{if $(i,j)=(2,2)$},\\ -4,&\textrm{if $(i,j)=(1,3)$ or $(3,1)$},\\ -6,&\textrm{if $(i,j)=(3,3)$},\end{array}\right.$
$\delta(m,i,n,j)=\left\{\begin{array}[]{ll}1,&\textrm{if $(m,i,n,j)\in G$ or $(n,j,m,i)\in G$},\\ 0,&\textrm{else},\end{array}\right.$
@@
We prove this lemma by mathematical induction. Firstly, we consider the case of $\alpha=1$. Since $l_{n,j}^{(1)}=0$ except $l_{1,0}^{(1)}=-4$ and $l_{1,2}^{(1)}=1$, from equation (4.11) we have
$\pi(D^{(1)}(l_{m,i},l_{n,j}))=\infty$
when $m,n\textgreater 1$,
$\pi(D^{(1)}(l_{1,i},l_{n,j}))=\pi(-D^{(1)}(l_{n,j},l_{1,i}))=\pi(l_{1,i}^{(1)}l_{n,j}^{(3)})\geq n+1+\lambda_{i,j}$
when $n\textgreater 1$, and
$\pi(D^{(1)}(l_{1,0},l_{1,2}))\geq 0=2+\lambda_{0,2},$
which prove (4.14) for $\alpha=1$. Specially, we obtain that
$\pi(D^{(1)}(l_{1,2},l_{2,0}))=\pi(l_{2,0}^{(3)})\geq 2,$
$\pi(D^{(1)}(l_{1,2},2l_{2,2}+l_{3,2}))=\pi(2l_{2,2}^{(3)}+l_{3,2}^{(3)})\geq 2,$
by Lemma 4.5, which prove that (4.15) and (4.16) hold for $\alpha=1$.
Secondly, we assume that (4.14)–(4.16) hold for $\alpha=s$ and prove that (4.17)–(4.19) hold for $\alpha=s$. We now compare the coefficients on both sides of
@@
Denote $p_{2}=y_{0}$ and $p_{3}=1$. For $u=0$ or $2$, define $x_{k,u}(m,i):=[p_{i}t^{m}]U^{(1)}(p_{u}t^{k})$ be the coefficient of $p_{i}t^{m}$ in $U^{(1)}(p_{u}t^{k})$, namely
@@
Then we derive that $l_{n,i}^{(2s)}=\sum\limits_{k,u}x_{k,u}(n,i)l_{k,u}^{(2s-1)}$ and
@@
By (4.7) and (4.8), we get $\pi(x_{k,u}(m,i))\geq\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor$, where $\gamma_{0,1}=1$, $\gamma_{2,1}=0$, $\gamma_{0,3}=-3$, and $\gamma_{2,3}=-4$. Now we first consider $(k,u,r,v)\in\{(1,2,2,2),(1,2,3,2),(3,2,1,2),(2,2,1,2)\}$. After checking all coefficients of $U^{(1)}(y_{0}t)$ and $U^{(1)}(y_{0}t^{2})$ given in the Appendix and $U^{(1)}(y_{0}t^{3})$ in (3.8), we obtain that
(4.21) $\pi(\sum_{r=2,3}{(x_{1,2}(m,i)x_{r,2}(n,j)+x_{r,2}(m,i)x_{1,2}(n,j))D^{(2s-1)}(l_{1,2},l_{r,2})})$
$\geq$ $2s-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
which are based on inequalities (4.14) and (4.16) when $\alpha=s$.

### v2

@@ 4. Radu and Sellers’ conjecture
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose that for $k=l$ or $k=l+1$ there exist polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$ such that
@@
(4.2) $\operatorname{ord}_{t}\!\left(p_{k}^{(i)}(t)\right)\geq\left\lceil\frac{k+s_{i}}{2}\right\rceil,\qquad i\in\{1,2\},$
for some fixed integers $s_{1}$ and $s_{2}$. Then there exist families of polynomials $p_{k}^{(1)}(t),p_{k}^{(2)}(t)\in\mathbb{Z}[t]$ for all integers $k\geq l$ such that (4.1) and (4.2) hold.
@@
Let $N>l+1$ be an integer and assume by induction that there exist families of polynomials $p_{k}^{(i)}(t)$, $i\in\{1,2\}$, such that (4.1) and (4.2) hold for $l\leq k\leq N-1$. Suppose
$p_{k}^{(i)}(t)=\sum_{n\geq\left\lceil\frac{k+s_{i}}{2}\right\rceil}c_{i}(k,n)\,t^{n},\qquad l\leq k\leq N-1,$
with integers $c_{i}(k,n)$. Applying Lemma 3.3 we obtain:
@@
$=-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n\geq\left\lceil\frac{N+j-2+s_{i}}{2}\right\rceil}c_{i}(N+j-2,n)\,t^{n}$
$=-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}a_{j}(t)t^{-1}\sum_{n\geq\left\lceil\frac{N+j+s_{i}}{2}\right\rceil}c_{i}(N+j-2,n-1)\,t^{n}.$
@@
Let $v_{1},v_{2},u:\mathbb{H}\rightarrow\mathbb{C}$ and $l\in\mathbb{Z}$. Suppose that for $k=l$ or $k=l+1$ there exist polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$, $i\in\{1,2\}$, such that
@@
(4.4) $p_{k}^{(i)}(t)=\sum_{n}c_{i}(k,n)\,2^{\left\lfloor\frac{2n-k+\gamma_{i}}{2}\right\rfloor}t^{n},$
with integers $\gamma_{i}$ and $c_{i}(k,n)$. Then there exist families of polynomials $p_{k}^{(i)}(t)\in\mathbb{Z}[t]$ for all integers $k\geq l$, of the form (4.4), for which property (4.3) holds.
@@
Suppose that for an integer $N>l+1$ there exist families of polynomials $p_{k}^{(i)}(t)$, $i\in\{1,2\}$, of the form (4.4) satisfying property (4.3) for $l\leq k\leq N-1$. We proceed by mathematical induction on $N$. Applying Lemma 3.3 and using the induction base (4.3) and (4.4) we obtain:
$U_{2}(ut^{N})=-\sum_{j=0}^{1}a_{j}(t)\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)\,2^{\left\lfloor\frac{2n-(N+j-2)+\gamma_{i}}{2}\right\rfloor}t^{n}.$
Utilizing (3.7) this rewrites as:
$U_{2}(ut^{N})$ $=-\sum_{j=0}^{1}\sum_{l=1}^{2}s(j,l)\,2^{\left\lfloor\frac{2l+j-1}{2}\right\rfloor}t^{l}\sum_{i=1}^{2}v_{i}\sum_{n}c_{i}(N+j-2,n)\,2^{\left\lfloor\frac{2n-(N+j-2)+\gamma_{i}}{2}\right\rfloor}t^{n}$
$=-\sum_{i=1}^{2}v_{i}\sum_{j=0}^{1}\sum_{l=1}^{2}\sum_{n}s(j,l)\,c_{i}(N+j-2,n-l)\,2^{\left\lfloor\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}\right\rfloor+\left\lfloor\frac{2l+j-1}{2}\right\rfloor}t^{n}.$
@@
$\left\lfloor\frac{2(n-l)-(N+j-2)+\gamma_{i}}{2}\right\rfloor+\left\lfloor\frac{2l+j-1}{2}\right\rfloor$
$\qquad\geq\left\lfloor\frac{2(n-l)-(N+j-2)+\gamma_{i}+2l+j-2}{2}\right\rfloor$
$\qquad=\left\lfloor\frac{2n-N+\gamma_{i}}{2}\right\rfloor.$
@@
A map $a:\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete function if it has finite support. A map $a:\mathbb{Z}\times\mathbb{Z}\rightarrow\mathbb{Z}$ is called a discrete array if for each $i\in\mathbb{Z}$ the map $a(i,-):\mathbb{Z}\rightarrow\mathbb{Z}$, $j\mapsto a(i,j)$, has finite support.
@@
Given $U^{(0)}(f)$ and $U^{(1)}(f)$ as in Definition 2.2, and $p_{0},p_{1},y_{0}$ as in Definition 2.5. Then there exist discrete arrays $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$, with $i\in\{0,1\}$, such that the following relations hold for all $k\in\mathbb{N}^{\ast}$:
(4.5) $U^{(0)}(t^{k})=p_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{0}(k,n)\,2^{\left\lfloor\frac{2n-k+2}{2}\right\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k-1}{2}\rceil}a_{1}(k,n)\,2^{\left\lfloor\frac{2n-k+2}{2}\right\rfloor}t^{n},$
(4.6) $U^{(0)}(p_{1}t^{k})=p_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{0}(k,n)\,2^{\left\lfloor\frac{2n-k+1}{2}\right\rfloor}t^{n}+y_{0}\sum_{n\geq\lceil\frac{k+1}{2}\rceil}b_{1}(k,n)\,2^{\left\lfloor\frac{2n-k-1}{2}\right\rfloor}t^{n},$
(4.7) $U^{(1)}(p_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}c_{0}(k,n)\,2^{\left\lfloor\frac{2n-k+1}{2}\right\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}c_{1}(k,n)\,2^{\left\lfloor\frac{2n-k-3}{2}\right\rfloor}t^{n},$
(4.8) $U^{(1)}(y_{0}t^{k})=p_{1}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{0}(k,n)\,2^{\left\lfloor\frac{2n-k}{2}\right\rfloor}t^{n}+\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{1}(k,n)\,2^{\left\lfloor\frac{2n-k-4}{2}\right\rfloor}t^{n}.$
@@
The Appendix lists eight fundamental relations. The two fundamental relations of Group I fit the pattern of relation (4.5) for two values of $k$. The same observation applies to Groups II, III, and IV with regard to relations (4.6), (4.8), and (4.7), respectively. Applying Lemmas 4.1 and 4.2, we immediately prove the statement for all $k\in\mathbb{N}^{\ast}$. ∎
@@
using Garvan’s MAPLE package ETA (3.1). We then note that $p$ and $p_{1}$ have expansions in powers of $q$ with coefficients in $\mathbb{Z}$.
Since $(1-q^{k})^{2}\equiv 1-q^{2k}\pmod{2}$, we have $\eta_{k}^{2}\equiv\eta_{2k}\pmod{2}$. Then
@@
which implies that the function $p$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$.
We now have that the coefficients of $4p_{1}=p+2pt+t+4t^{2}$ are integral. Hence, by the relations for $U^{(1)}(p_{0}t^{2})$ given in the Appendix,
@@
where $f(t)$ and $g(t)$ are polynomials in $t$ with integral coefficients, we see that $p_{1}t$ has expansions in powers of $q$ with coefficients in $\mathbb{Z}$ as does $p_{1}$. Furthermore, using (4.9) and combining the relations for $U^{(1)}(p_{0}t)$, $U^{(1)}(y_{0}t)$, $U^{(0)}(p_{1}t)$ given in the Appendix together with $U^{(0)}(t^{4})$ given in (3), we now state the following representation.
@@
$L_{3}$ $=y_{0}t+2y_{0}t^{2}+4g(p_{0},y_{0},t),$
for some polynomials $f$ and $g$ with integral coefficients.
@@
For $\alpha\geq 1$ there exist integers $d_{n,i}^{(\alpha)}$, $i=0,1,2,3$, such that
@@
By (4.9) we see that $L_{1}=-4p_{0}t+y_{0}t$ has the form given in (4.11). Assume that $L_{2\alpha-1}$ has the form given in (4.11) for a fixed $\alpha$. Then
@@
We aim to show that each sum on the right-hand side of (4.13) satisfies the form in (4.12). From (4.7) and (4.8),
$L_{2\alpha}$ $=p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{0}(k,n)\,2^{k-1+\left\lfloor\frac{2n-k+1}{2}\right\rfloor}t^{n}$
$\quad+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,0}^{(2\alpha-1)}c_{1}(k,n)\,2^{k-1+\left\lfloor\frac{2n-k-3}{2}\right\rfloor}t^{n}$
$\quad+p_{1}\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{0}(k,n)\,2^{k-1+\left\lfloor\frac{2n-k}{2}\right\rfloor}t^{n}$
$\quad+\sum_{k=1}^{\infty}\sum_{n\geq\lceil\frac{k+4}{2}\rceil}d_{k,2}^{(2\alpha-1)}d_{1}(k,n)\,2^{k-1+\left\lfloor\frac{2n-k-4}{2}\right\rfloor}t^{n}.$
@@
$k-1+\left\lfloor\frac{2n-k+1}{2}\right\rfloor$ $=\left\lfloor\frac{2n-k+1+2k-2}{2}\right\rfloor=n+\left\lfloor\frac{k-1}{2}\right\rfloor\geq n,$
$k-1+\left\lfloor\frac{2n-k}{2}\right\rfloor$ $=\left\lfloor\frac{2n-k+2k-2}{2}\right\rfloor=n+\left\lfloor\frac{k-2}{2}\right\rfloor\geq n-1.$
@@
$k-1+\left\lfloor\frac{2n-k-3}{2}\right\rfloor$ $=\left\lfloor\frac{2n-k-3+2k-2}{2}\right\rfloor=n+\left\lfloor\frac{k-5}{2}\right\rfloor\geq n-2,$
$k-1+\left\lfloor\frac{2n-k-4}{2}\right\rfloor$ $=\left\lfloor\frac{2n-k-4+2k-2}{2}\right\rfloor=n+\left\lfloor\frac{k-6}{2}\right\rfloor\geq n-3.$
Thus the sums on the right-hand side of (4.13) have the form of (4.12). Hence $L_{2\alpha}$ has the desired form. The proof that the correct form of $L_{2\alpha}$ implies the correct form of $L_{2\alpha+1}$ is analogous. The general result follows by induction. ∎
By Lemma 4.2 and Lemma 4.6 we set
@@
and denote by $\pi(n)$ the $2$-adic order of $n$ (i.e., the highest power of $2$ that divides $n$). By Lemma 4.6, we note that for $m,n\in\mathbb{N}$ and $\alpha\geq 1$, the coefficients $l_{n,1}$ and $l_{m,3}$ appear only in $L_{2\alpha}$, while $l_{n,0}$ and $l_{m,2}$ appear only in $L_{2\alpha-1}$.
@@
    (4.14) $\pi\bigl(D^{(2\alpha-1)}(l_{m,i},l_{n,j})\bigr)\geq 2\alpha-2+m+n+\lambda_{i,j},$
    (4.15) $\pi\bigl(D^{(2\alpha-1)}(l_{1,2},l_{2,0})\bigr)\geq 2\alpha,$
    (4.16) $\pi\bigl(D^{(2\alpha-1)}(l_{1,2},2l_{2,2}+l_{3,2})\bigr)\geq 2\alpha;$
@@
    (4.17) $\pi\bigl(D^{(2\alpha)}(l_{m,i},l_{n,j})\bigr)\geq 2\alpha-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
    (4.18) $\pi\bigl(D^{(2\alpha)}(l_{1,1},l_{3,3})\bigr)\geq 2\alpha,$
    (4.19) $\pi\bigl(D^{(2\alpha)}(l_{2,1},l_{3,3})\bigr)\geq 2\alpha+1;$
@@
[…]
$\delta(m,i,n,j)=\left\{\begin{array}[]{ll}1,&\text{if }(m,i,n,j)\in G\text{ or }(n,j,m,i)\in G,\\ 0,&\text{otherwise},\end{array}\right$.
@@
We prove this lemma by mathematical induction. First, we consider the case $\alpha=1$. Since $l_{n,j}^{(1)}=0$ except $l_{1,0}^{(1)}=-4$ and $l_{1,2}^{(1)}=1$, from equation (4.11) we have
$\pi\bigl(D^{(1)}(l_{m,i},l_{n,j})\bigr)=\infty$
when $m,n>1$,
$\pi\bigl(D^{(1)}(l_{1,i},l_{n,j})\bigr)=\pi\bigl(-D^{(1)}(l_{n,j},l_{1,i})\bigr)=\pi\bigl(l_{1,i}^{(1)}l_{n,j}^{(3)}\bigr)\geq n+1+\lambda_{i,j}$
when $n>1$, and
$\pi\bigl(D^{(1)}(l_{1,0},l_{1,2})\bigr)\geq 0=2+\lambda_{0,2},$
which proves (4.14) for $\alpha=1$. In particular, we obtain that
$\pi\bigl(D^{(1)}(l_{1,2},l_{2,0})\bigr)=\pi\bigl(l_{2,0}^{(3)}\bigr)\geq 2,$
$\pi\bigl(D^{(1)}(l_{1,2},2l_{2,2}+l_{3,2})\bigr)=\pi\bigl(2l_{2,2}^{(3)}+l_{3,2}^{(3)}\bigr)\geq 2,$
by Lemma 4.5, which proves that (4.15) and (4.16) hold for $\alpha=1$.
Second, we assume that (4.14)–(4.16) hold for $\alpha=s$ and prove that (4.17)–(4.19) hold for $\alpha=s$. We now compare the coefficients on both sides of
@@
Denote $p_{2}=y_{0}$ and $p_{3}=1$. For $u=0$ or $2$, define $x_{k,u}(m,i):=[p_{i}t^{m}]U^{(1)}(p_{u}t^{k})$ to be the coefficient of $p_{i}t^{m}$ in $U^{(1)}(p_{u}t^{k})$, namely
@@
Then we derive that $l_{n,i}^{(2s)}=\sum_{k,u}x_{k,u}(n,i)l_{k,u}^{(2s-1)}$ and
@@
By (4.7) and (4.8), we obtain $\pi\bigl(x_{k,u}(m,i)\bigr)\geq\left\lfloor\frac{2m-k+\gamma_{u,i}}{2}\right\rfloor$, where $\gamma_{0,1}=1$, $\gamma_{2,1}=0$, $\gamma_{0,3}=-3$, and $\gamma_{2,3}=-4$. Now we first consider $(k,u,r,v)\in\{(1,2,2,2),(1,2,3,2),(3,2,1,2),(2,2,1,2)\}$. After checking all coefficients of $U^{(1)}(y_{0}t)$ and $U^{(1)}(y_{0}t^{2})$ given in the Appendix and $U^{(1)}(y_{0}t^{3})$ in (3), we obtain that
$\pi\!\left(\sum_{r=2,3}\bigl(x_{1,2}(m,i)x_{r,2}(n,j)+x_{r,2}(m,i)x_{1,2}(n,j)\bigr)D^{(2s-1)}(l_{1,2},l_{r,2})\right)$
(4.21) $\qquad\geq 2s-1+m+n+\lambda_{i,j}-\delta(m,i,n,j),$
which is based on inequalities (4.14) and (4.16) for $\alpha=s$.